Assume that in a particular year, the natural rate of unemployment is 5 percent and the actual rate of unemployment is 9 percent. Use the Okun’s law to estimate the size of GNP gap in percentage point terms. If the potential GNP is 500 billion in that year, how much output is being foregone because of cyclical unemployment?


1
Expert's answer
2020-08-19T14:21:01-0400

Solution:

Okun's law states that: For every 1% increase in cyclical unemployment (actual unemployment – natural rate of unemployment), GDP will decrease by -2%.

Therefore:

a).

"GNP \; Gap=(Actual \; rate\; of\; unemployment-Natural\; rate\; of\; unemployment) \times-2\%"


"=(9\%-5\%) \times-2\%"

"=(4\%) \times -2\%"


"=-8\%"


The size of GNP Gap in percentage point terms = "=-8\%"


b). The output that is being foregone because of cyclical unemployment if the potential GNP is 500 billion, will be:

"=Potential \;GNP\times GNP\;Gap \;in \;percentage\;point\;terms"

"=\$500 \;billion\times -8\%"

"=\$40\;billion"